Calculate the heat change in kilojoules for vaporization of 4.00 kg of water at 100 C

Please express answer as a positive value using three significant figures and include the appropriate units.
1
Expert's answer
2020-12-18T08:36:41-0500

Change in gases moles = 4000/18 = 222.23

Applying

Change in heat = (change in number of moles)(gas constant)(Temperature)

Change of heat = 222.23 (8.314)(373)/1000

= 689.16 kilojoule/mole